solve using quadratic equation
4x²+7x-15=0

Answer by stanbon(75887)

solve using quadratic equation
4x²+7x-15=0
x = [-7 +- sqrt(49-4*4*-15)]/8
x = [ -7 +- sqrt289)]/8
x = [-7 +- 17]/8
x = -24/8 or x = 10/8
x = -3 or x = 5/4
====================
Cheers,
Stan H.